OpenStudy (anonymous):

How would you write twenty-five squared divided by five to the fourth power, and evaluate it?

OpenStudy (anonymous):

Well you would write it as: \[\frac{ 25^2 }{ 5^4 }\] And eval that as: \[\frac{ 25^2 }{ 5^4 } = \frac{ 625 }{ 625 } = 1\]